OpenStudy (anonymous):

how is the following solved? b/b^2-25 +5/b+5-6/b

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

It's not an equation...

OpenStudy (anonymous):

do you want to simplify it?

OpenStudy (anonymous):

oh you are trying to combine them into one fraction yes?

OpenStudy (anonymous):

then the common denominator is b(b+5)(b-5). get everything in that form and combine

OpenStudy (anonymous):

\[\frac{b}{b^2-25}+\frac{5}{b+5}-\frac{6}{b}\]

OpenStudy (anonymous):

multiply first one top and bottom by b multiply second one by \[(b-5)b\] and third by \[(b+5)(b-5)\] then put them over the same denominator and multiply out in the numeator

OpenStudy (anonymous):

\[\frac{b\times b + 5\times b(b-5)-6(b+5)(b-5)}{(b-5)(b+5)b}\]

OpenStudy (anonymous):

leave the denominator alone, multiply out in the numerator \[b^2+5b^2-25b-6b^2-150\] \[-25b-150\] is your numerator if i did the algebra correctly

OpenStudy (anonymous):

nope i made a mistake!

OpenStudy (anonymous):

last term has -6*-25 = 150 not -150 sorry. answer is \[\frac{-25b+150}{b(b+5)(b-5)}\]